Given that Z is a standard normal random variable, find Z for each

situation.

a) The area between -Z and Z is .9030.

b) The area between -Z and Z is .9528.


1
Expert's answer
2021-07-14T13:01:22-0400

a)

"P(-z < Z < z) = 0.9030 \\

P(Z < z) - (1-P(Z < z)) = 0.9030 \\

2P(Z < z) - 1 = 0.9030 \\

2P(Z < z) = 1 + 0.9030 \\

2P(Z < z) = 1.9030 \\

P(Z < z) = \frac{1.9030 }{ 2 } \\

P(Z < z) = 0.9515"

See the probability 0.9515 in the standard normal table the corresponding z value 1.66

"P(Z < 1.66) = 0.9515 \\

z = 1.66"

b)

"P(-z < Z < z) = 0.9528 \\

P(Z < z) - (1-P(Z < z)) = 0.9528 \\

2P(Z < z) - 1 = 0.9528 \\

2P(Z < z) = 1 + 0.9528 \\

2P(Z < z) = 1.9528 \\

P(Z < z) = \frac{1.9528 }{ 2 } \\

P(Z < z) = 0.9764"

See the probability 0.9764 in the standard normal table the corresponding z value 1.98

"P(Z< 1.98) = 0.9764 \\

z= 1.98"